OpenStudy (anonymous):

how many atoms of C are there in 25 grams of C5H12 ? (pentane)

OpenStudy (anonymous):

LOOK AT the molecular formula of pentane .82 grams is molecular mass of pentane which is also molar mass of pentane now 1 mole of pentane contains 5*AVAGADROS NUMBER carbons atoms .now calculate the number of moles in 25 grams i hope u illl understand

OpenStudy (xishem):

In a more mathematical sense:\[25g(C_5H_{12})*\frac{1mol(C_5H_{12})}{72.15g(C_5H_{12})}=0.346500mol(C_5H_{12})\]\[0.346500mol(C_5H_{12})*\frac{6.022*10^{23}mlcl(C_5H_{12})}{1mol(C_5H_{12})}=2.086625*10^{23}mlcl(C_5H_{12}))\] Since each molecule of pentane contains 5 atoms of carbon:\[2.086625*10^{23}mlcl(pentane)*\frac{5atoms(C)}{1mlcl(pentane)}=1.0*10^{24}atoms(C)\]We're only justified in keeping 2sigfigs in the final answer.